1. Morphine 4 mg IVP is ordered. Available is 10 mg/mL. You further dilute to a total of 10 mL and plan to administer over 5 min. After 3 minutes you notice your patient’s respirations have dropped below 10 and you stop the medication. How many mg were administered?
Answer #1

Ans) 1) Order: Morphine 4 mg IVP

Available: Morphine 10 mg/ml

Dilute further: 100 mg = 10 ml

4 mg = x

x = 40/100 = 0.4 ml

To infuse in 5 mins: 0.4/5 = 0.08 ml/min

In 3 mins: 0.08 × 3 = 0.24 ml.